a set of data is distributed with a mean of 98 and a standard deviation of 5. What is the probability that a value from the data is less than 88?
Answer by stanbon(75887)

a set of data is distributed with a mean of 98 and a standard deviation of 5. What is the probability that a value from the data is less than 88?
-----
z(88) = (88-98)/5 = --2
---
P(x < 88) = P(z < -2) = 0.0228
=======================
Cheers,
Stan H.